please explain step process in detail

On January 2, 2016, Sweet Pet purchased fixtures for $54,100 cash, expecting the fxtures to remain in service for seven years. Sweet Pet has depreciated the fixtures on a straight-line basis, with $10,000 residual value. On May 31, 2018, Sweet Pet sold the fixtures for $35,375 cash. Record both depreciation expense for 2018 and sale of the fixtures on May 31, 2018. (Assume the modified half-month convention is used. Record debits first, then credits. Select the explanation on the last line of the journal entry table.) Begin by recording the depreciation expense as of May 31, 2018 Date Accounts and Explanation Debit Credit May 31
Answer #1

Workings:

Calculation of Depreciation Expense:

Straight line depreciation expense = (Cost of fixtures - residual value)/useful life

=($54100 - $10000)/7 = $44100 / 7 = $6300

Straight line depreciation expense for year 2018 = $6300*5/12 = $2625

Accumulated depreciation up to 31 may 2018 = $6300 + $6300 + $2625 = $15225

Book value of fixtures as on 31 may 2018 = Cost of fixtures - Accumulated depreciation = $54100 - $15225 = $38875

Sale price of fixtures = $35375

Loss on sale of Fixtures = Book value of fixtures - Sale price of fixtures = $38875 - $35375 = $3500

Journal entries:

Date

Accounts and Explanation

Debit($)

Credit($)

May 31,2018

Depreciation Expense

2625

Accumulated Depreciation

2625

(Depreciation recorded as on the date of sale)

May 31,2018

Cash

35375

Accumulated Depreciation

15225

Loss on sale of Fixtures

3500

   Fixtures

54100

(Sale of fixtures recorded)